Taffer Consuluting Services would like to go public to raise $90 million to support expected growth. Their investment bank charges the following:

  • 6.6% underwriting spread for a firm commitment
  • $421,907 in legal fees

The underwriter feels that the IPO will be priced at $5.09 per share. How many shares must be issued to net the company the proceeds it needs to fund its future investment (Hint: The total amount issued must cover the required net proceeds plus all fees and the underwriter's spread)?
Answer #1

Total amount to be raised
=(Amount Received+Administrative Costs)/(1- Spread)= (90,000,000+421907)/(1-6.6%) =96811463.5974

Number of shares =Total Amount to be received/Price per share =96811463.5974/5.09 =19,019,933.91 shares